# SVIP: Towards Verifiable Inference of Open-source Large Language Models

### v1

@@ Abstract
Open-source Large Language Models (LLMs) have recently demonstrated remarkable capabilities in natural language understanding and generation, leading to widespread adoption across various domains. However, their increasing model sizes render local deployment impractical for individual users, pushing many to rely on computing service providers for inference through a blackbox API. This reliance introduces a new risk: a computing provider may stealthily substitute the requested LLM with a smaller, less capable model without consent from users, thereby delivering inferior outputs while benefiting from cost savings. In this paper, we formalize the problem of verifiable inference for LLMs. Existing verifiable computing solutions based on cryptographic or game-theoretic techniques are either computationally uneconomical or rest on strong assumptions. We introduce SVIP, a secret-based verifiable LLM inference protocol that leverages intermediate outputs from LLM as unique model identifiers. By training a proxy task on these outputs and requiring the computing provider to return both the generated text and the processed intermediate outputs, users can reliably verify whether the computing provider is acting honestly. In addition, the integration of a secret mechanism further enhances the security of our protocol. We thoroughly analyze our protocol under multiple strong and adaptive adversarial scenarios. Our extensive experiments demonstrate that SVIP is accurate, generalizable, computationally efficient, and resistant to various attacks. Notably, SVIP achieves false negative rates below $5\%$ and false positive rates below $3\%$, while requiring less than $0.01$ seconds per query for verification.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have achieved unprecedented success across a broad array of tasks and domains (Achiam et al., 2023; Dubey et al., 2024; Yang et al., 2024). Alongside this progress, open-source LLMs have proliferated, offering increasingly sophisticated and capable models to the broader research community (Touvron et al., 2023b; Black et al., 2022; Le Scao et al., 2023; Jiang et al., 2023; Almazrouei et al., 2023; Zhang et al., 2023). Many of these open-source LLMs now rival, or even surpass, their closed-source counterparts in performance (Chiang et al., 2023; Almazrouei et al., 2023; Dubey et al., 2024), while remaining freely accessible. However, as model capacity grows, it typically comes with a corresponding increase in the number of model parameters, which directly drives up the computational demands, particularly in terms of memory and processing power (Kukreja et al., 2024). For example, the Llama-3.1-70B model requires $140$ GB of GPU memory for inference under FP16 (Hugging Face, 2024). These escalating computational costs have made local deployment increasingly impractical, pushing many users to turn to computing service providers for inference.
@@
To ease the deployment of LLM inference, computing providers may provide API-only access to users, hiding implementation details. A new risk arises in this setting: how to ensure that the outputs from a computing provider are indeed generated by the requested LLM? For instance, a user might request the Llama-3.1-70B model for complex tasks, but a dishonest computing provider could substitute the smaller Llama-2-7B model for cost savings, while still charging for the larger model. The smaller model demands significantly less memory and processing power, giving the computing provider a strong incentive to cheat. Restricted by the black-box API access, it is difficult for the user to detect model substitution. Particularly, future computing service may be provided by decentralized individuals (Uriarte & DeNicola, 2018) who sell their computation power at a competitive price, but the computation outputs may not be trustworthy. This highlights the need for verifiable inference, a mechanism designed to ensure that the model specified by the user is the one actually used during inference. Implementing verifiable inference for LLMs is essential not only for safeguarding users’ interests but also for fostering trust in the open-source LLM ecosystem, driving low-cost and wider adoption and continued development of more advanced LLMs.
@@
Figure 1: The problem setting of verifiable inference for LLMs. (a) A user requests the computing provider (referred to as provider in the figure) to run inference on their prompt using the Llama-3.1-70B model. Without verification, they have no way to confirm if the specified model is used. (b) Our proposed protocol solves this by requiring the provider to return processed intermediate outputs from the LLM, enabling the user to verify through a verification function whether the correct model was used for inference. Specifically, the intermediate outputs are compressed to reduce the computational overhead.
@@
An effective verifiable inference solution for LLMs must accurately confirm that the specified model is being used during inference, while maintaining computational efficiency. Simple solutions, such as probing the model with established benchmark datasets, may be easily detected and bypassed. On the other hand, cryptographic verifiable computing methods, which rely on generating mathematical proofs (Yu et al., 2017; Setty et al., 2012) or secure computation techniques (Gennaro et al., 2010; Laud & Pankova, 2014), are often too computationally expensive for real-time LLM inference. For instance, zkLLM, a recent Zero Knowledge Proof-based technique, requires over $803$ seconds for a single prompt query (Sun et al., 2024). Game-theoretic protocols such as Zhang et al. (2024) involve the interaction of multiple computing providers with carefully designed penalties and rewards, assuming all providers are rational, flawless, and non-cooperative, which might be unrealistic in practice. Meanwhile, watermarking and fingerprinting techniques (Kirchenbauer et al., 2023; Xu et al., 2024) are mostly implemented by model publishers, making them unsuitable for verifiable inference, where the verification primarily occurs between the user and the computing provider.
@@
In this paper, we propose SVIP, a Secret-based Verifiable LLM Inference Protocol using intermediate outputs. The core idea of our method is to require the computing provider to return not only the generated text but also the processed intermediate outputs (hidden state representations) from the LLM. We carefully design and train a proxy task exclusively on the hidden representations produced by the specified model, effectively transforming these representations into a distinct identifier for that model. During deployment, users can verify whether the processed hidden states returned by the computing provider come from the specified model by assessing their performance on the proxy task. If the returned outputs perform well on this task, it provides strong evidence that the correct model was used for inference. We further strengthen the security of our protocol with a secret-based mechanism, making it difficult for a malicious computing provider to fake or bypass the verification process. We also conduct a thorough security analysis, addressing potential attacks such as direct vector optimization attack, adapter attack, and secret recovery attack. Our key contributions are:
  • We take the first step toward systematically formalizing the problem of verifiable LLM inference and propose an innovative protocol that leverages processed intermediate outputs. Notably, our protocol does not require retraining or fine-tuning the LLM. The security of our protocol is further enhanced by a novel secret-based mechanism.
@@
  • Our comprehensive experiments with $5$ specified open-source LLMs (ranging from $13$B to $70$B) demonstrate the effectiveness of SVIP: it achieves an average false negative rate of $3.49\%$, while maintaining the false positive rate below $3\%$ across $6$ smaller alternative models. SVIP introduces negligible overhead (less than $0.01$ seconds per query) for both the user and the computing provider.
  • We provide a thorough discussion and analysis of various strong and adaptive attack scenarios in verifiable LLM inference. Our results show that SVIP can effectively and securely handle approximately $80$ to $120$ million prompt queries in total after a single round of protocol training, with the update mechanism further bolstering security.
@@ 2 Problem Statement
We begin by formalizing the verifiable inference problem in the context of LLMs. We consider three parties: the user, the computing provider, and a trusted third party. The user intends to use a specified LLM, $\mathcal{M}_{spec}$, to perform inference on a prompt $x\in\mathcal{V}^{*}$, where $\mathcal{V}^{*}$ represents the set of all possible string sequences for a vocabulary set $\mathcal{V}$. Lacking sufficient computational resources to run $\mathcal{M}_{spec}$ locally, the user relies on the computing provider to execute the model and accordingly pays for the service. Ideally, the computing provider would run $\mathcal{M}_{spec}$ as requested and return the completion. However, a dishonest provider might stealthily substitute an alternative LLM, $\mathcal{M}_{alt}$, which could be significantly smaller than $\mathcal{M}_{spec}$ in terms of model size, and return an inferior result.
The goal of a trusted third party, is to design and implement a verification protocol that verifies whether the computing provider uses $\mathcal{M}_{spec}$ for inference. Based on this protocol, the user can determine with high confidence whether the computing provider used $\mathcal{M}_{spec}$ (True) or not (False) for any query $x$. A satisfactory protocol should meet the following criteria:
@@
    Low False Negative Rate (FNR): The protocol should minimize instances where the computing provider did use $\mathcal{M}_{spec}$ but is wrongly flagged as not using it.
@@
    Low False Positive Rate (FPR): The protocol should minimize cases where it incorrectly confirms that the computing provider used $\mathcal{M}_{spec}$ when, in fact, it used an alternative model, $\mathcal{M}_{alt}$.
@@
    Efficiency: The verification protocol should be computationally efficient and introduce minimal overhead for both the computing provider and the user.
  4. 4.
    Preservation of Completion Quality: The protocol should not compromise the quality of the prompt completion returned by the computing provider.
3 Related work
Verifiable Computing
Verifiable Computing (VC) allows users to verify that an untrusted computing provider has executed computations correctly, without having to perform the computation themselves (Walfish & Blumberg, 2015; Yu et al., 2017; Costello et al., 2015; Kosba et al., 2018). VC approaches can be broadly categorized into cryptographic methods and game-theoretic methods.
Cryptographic VC techniques either require the provider to return a mathematical proof that confirms the correctness of the results (Ghodsi et al., 2017; Setty et al., 2012; Parno et al., 2016), or rely on secure computation techniques (Gennaro et al., 2010; Madi et al., 2020; Laud & Pankova, 2014). These techniques cryptographically guarantee correctness and have been applied to machine learning models and shallow neural networks (Niu et al., 2020; Zhao et al., 2021; Hu et al., 2023a; Lee et al., 2024; Ghodsi et al., 2017; Lee et al., 2022). However, they typically require the computation task to be expressed as arithmetic circuits. Representing open-source LLMs in circuit form is particularly challenging due to their complex architectures and intricate operations. Moreover, the sheer size of these models, with billions of parameters, introduces substantial computational overhead. A recent work, zkLLM (Sun et al., 2024), attempts to verify LLM inference using Zero Knowledge Proofs. For the Llama-2-13B (Touvron et al., 2023b) model, generating a proof for a single prompt takes $803$ seconds, and repeating this process for large batches of prompt queries becomes computationally prohibitive.
In contrast, game-theoretic VC techniques ensure the correctness of outsourced computations by leveraging economic incentives to enforce honest behavior (Nabi et al., 2020; Liu & Zhang, 2018). For instance, a sampling-based verification mechanism Proof of Sampling (Zhang et al., 2024) requires multiple computing service providers to independently compute and compare results, ensuring integrity through penalties and rewards. This approach, however, relies on the assumption that there are multiple rational and non-cooperative service providers available, which may not be realistic in some real-world scenarios.
LLM Watermarking and Fingerprinting
LLM watermarking involves embedding algorithmically detectable signals into the text generated by LLMs, with the goal of identifying AI-generated texts (Kirchenbauer et al., 2023; Hu et al., 2023b; Christ et al., 2024; Gu et al., 2023). Meanwhile, LLM fingerprinting implants specific backdoor triggers into LLMs, causing the model to generate particular text whenever a confidential private key is used (Xu et al., 2024). Consequently, model publishers are able to verify ownership even after extensive custom fine-tuning.
@@ 4 Methodology
(a) Simple Protocol
(b) Secret-based Protocol
Figure 2: Illustration of (a) simple protocol (Section 4.1); (b) secret-based protocol (Section 4.2).
@@
Despite the fact that larger language models typically offer superior text generation quality (Kaplan et al., 2020), it is often challenging to verify whether a computing provider is using $\mathcal{M}_{spec}$ for inference based solely on the returned completion text. Our framework addresses this by requiring the computing provider to return not only the generated text but also the processed intermediate outputs (hidden state representations) from the LLM inference process.
We design and train a proxy task specifically to perform well only on the hidden representations generated by $\mathcal{M}_{spec}$ during the protocol’s training stage. The intuition behind is that the proxy task transforms the hidden representations into a unique identifier for the model. During deployment, the user can evaluate the performance of the returned intermediate outputs on the proxy task. Strong performance on the proxy task indicates that the correct model was used for inference, while poor performance suggests otherwise.
Our approach does not depend on expensive cryptographic proofs or protocols, and is highly efficient. Furthermore, it does not involve retraining or fine-tuning the LLMs, operates independently of the model publisher, and can be applied to any LLM with publicly available weight parameters, making it widely applicable. Next, we formalize and illustrate our proposed framework in detail.
4.1 A Simple Protocol Based on Intermediate Outputs
@@
We denote the prompt input as $x\in\mathcal{V}^{*}$. For any LLM $\mathcal{M}$, let $h_{\mathcal{M}}(x)\in\mathbb{R}^{L\times d_{\mathcal{M}}}$ represent the last-layer hidden representations of $x$ produced by $\mathcal{M}$, where $L$ is the length of the tokenized input $x$, and $d_{\mathcal{M}}$ denotes the hidden dimension of $\mathcal{M}$. The computing provider receives $x$ from the user, runs $\mathcal{M}$, and returns $h_{\mathcal{M}}(x)$ to user for subsequent verification. However, to reduce the size of the intermediate outputs returned, we additionally apply a proxy task feature extractor network $g_{\theta}(\cdot):\mathbb{R}^{L\times d_{\mathcal{M}}}\rightarrow\mathbb{R}^{d_{g}}$ parameterized by $\theta$, where $d_{g}$ represents the proxy task feature dimension. The computing provider now also runs $g_{\theta}(\cdot)$ and returns a compressed vector $z(x):=g_{\theta}(h_{\mathcal{M}}(x))$ of dimension $d_{g}$ to the user, significantly reducing the communication overhead. Specifically, for each prompt query, the compressed vector only takes approximately $4$ KB when $d_{g}$ is set to $1024$.
The user is required to perform two tasks locally: obtaining the predicted proxy task output and the label. First, the user runs $f_{\phi}(\cdot)$, using the returned proxy task feature $z(x)$ as input to compute $f_{\phi}(z(x))$. Here, $f_{\phi}(\cdot):\mathbb{R}^{d_{g}}\rightarrow\mathcal{Y}$ is the proxy task head parameterized by $\phi$, where $\mathcal{Y}$ denotes the label space. Second, the user applies a labeling function for the proxy task. We adopt a self-labeling function $y(x):\mathcal{V}^{*}\rightarrow\mathcal{Y}$, which derives the label directly from the input, eliminating the need for external labels or specialized annotators [^1] For instance, we can define $y(x)$ as the Set-of-Words (SoW) representation of the input $x$, which captures the presence of each word in a fixed vocabulary, regardless of frequency. As a concrete example, if $\mathcal{V}=\{a,b,c,d\}$ and $x=``abcc"$, the SoW label $y(x)$ would be a four-dimensional vector $(1,1,1,0)$, indicating whether each token in $\mathcal{V}$ appears in $x$.. Note that the label can either be a scalar or a vector.
@@
A trusted third party is responsible for implementing the protocol. With a properly defined loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ and a training dataset $\mathcal{D}$, the trusted third party trains the proxy task according to the following training objective:
@@
Low $\begin{split}\quad&\textbf{Low FNR}:\mathbb{P}\left(V(x,z(x);\phi^{*},\theta^{*})=0|\mathcal{M}_{spec}~{}\text{is used for inference}\right)\leq\alpha;\\ \quad&\textbf{Low FPR}:\mathbb{P}\left(V(x,z(x);\phi^{*},\theta^{*})=1|\mathcal{M}_{spec}~{}\text{is {not} used for inference}\right)\leq\beta.\end{split}$ (2)
While a single prompt query may occasionally yield an incorrect verification result due to FNR or FPR, in practice, users can perform the verification over a batch of queries and apply a hypothesis testing to reach a conclusion with high confidence. Refer to Appendix A.3 for a detailed discussion.

### v2

@@ Abstract
The ever-increasing size of open-source Large Language Models (LLMs) renders local deployment impractical for individual users. Decentralized computing has emerged as a cost-effective solution, allowing individuals and small companies to perform LLM inference for users using surplus computational power. However, a computing provider may stealthily substitute the requested LLM with a smaller, less capable model without consent from users, thereby benefiting from cost savings. We introduce SVIP, a secret-based verifiable LLM inference protocol. Unlike existing solutions based on cryptographic or game-theoretic techniques, our method is computationally effective and does not rest on strong assumptions. Our protocol requires the computing provider to return both the generated text and processed hidden representations from LLMs. We then train a proxy task on these representations, effectively transforming them into a unique model identifier. With our protocol, users can reliably verify whether the computing provider is acting honestly. A carefully integrated secret mechanism further strengthens its security. We thoroughly analyze our protocol under multiple strong and adaptive adversarial scenarios. Our extensive experiments demonstrate that SVIP is accurate, generalizable, computationally efficient, and resistant to various attacks. Notably, SVIP achieves false negative rates below $5\%$ and false positive rates below $3\%$, while requiring less than $0.01$ seconds per prompt query for verification.
@@ 1 Introduction
In recent years, open-source Large Language Models (LLMs) have achieved unprecedented success across a broad array of tasks and domains [44, 2, 26, 19], while remaining freely accessible. However, as model sizes increase, so do their computational demands [23]. As a result, decentralized computing [45] has gained significant attention as a cost-effective solution for users with limited local computational resources. In this setting, a user lacking computational power relies on decentralized computing providers to perform LLM inference. These providers, often individuals or small companies with surplus resources, offer computational power at competitive prices. Commercial platforms facilitate such interactions by connecting both parties. Real-world examples include Golem Network, Akash Network, Render Network, Spheron Network, Hyperbolic, and Vast.ai.
@@
However, unlike reputable companies with well-established credibility, computation outputs from decentralized computing providers may not always be trustworthy. Specifically, to ease the deployment of LLM inference, computing providers often provide API-only access to users, hiding implementation details. A new risk arises in this setting: how to ensure that the outputs from a computing provider are indeed generated by the requested LLM? For instance, a user might request the Llama-3.1-70B model for complex tasks, but a dishonest computing provider could substitute the smaller Llama-2-7B model for cost savings, while still charging for the larger model. The smaller model demands significantly less memory and processing power, giving the computing provider a strong incentive to cheat. Restricted by the black-box API access, it is difficult for the user to detect such substitutions.
Without assurance that users are receiving the service they specified and paid for, they may lose trust and abandon the platform. To prevent this outcome and maintain profitability, the platform must ensure that user-specified models are faithfully executed. This highlights the need for verifiable inference, a mechanism designed to ensure that the model specified by the user is the one actually used during inference.
@@
Figure 1: The problem setting of verifiable inference for LLMs. (a) Our protocol involves three parties. (b) A user requests the computing provider (referred to as provider in the figure) to run inference on their prompt using the Llama-3.1-70B model. Without verification, they have no way to confirm if the specified model is used. (c) Our proposed protocol solves this by requiring the provider to return processed hidden representations from the LLM, enabling the user to verify through a verification function whether the correct model was used for inference. Specifically, the hidden representations are compressed to reduce the computational overhead.
Related Work
@@
A practical verifiable inference solution for LLMs must accurately confirm that the specified model is being used during inference while maintaining computational efficiency. Existing approaches face significant challenges that limit their applicability. Cryptographic verifiable computing methods, which rely on generating mathematical proofs [50, 38] or secure computation techniques [13, 25] often impose high computational costs, making them unsuitable for real-time LLM inference. For instance, zkLLM, a recent Zero Knowledge Proof-based technique, requires over $803$ seconds for a single prompt query [40]. Game-theoretic protocols involve the interaction of multiple computing providers with carefully designed penalties and rewards [53], assuming all providers are rational, flawless, and non-cooperative, which might be unrealistic for certain system setups in practice. Meanwhile, watermarking and fingerprinting techniques [21, 48] are mostly implemented by model publishers, making them unsuitable for verifiable inference, where the verification primarily occurs between the user and the computing provider. We leave an extended discussion of related work to Appendix D.
@@
In this paper, we propose SVIP, a Secret-based Verifiable LLM Inference Protocol using hidden representations. Our protocol requires the computing provider to return not only the generated text but also the processed hidden state representations from the LLM. We carefully design and train a proxy task exclusively on the hidden representations produced by the specified model, effectively transforming these representations into a distinct identifier for that model. During deployment, users can verify whether the processed hidden states returned by the computing provider come from the specified model by assessing their performance on the proxy task. If the returned representations perform well on this task, it provides strong evidence that the correct model was used for inference. Our key contributions are:
  • We systematically formalize the problem of verifiable LLM inference (§2) and propose an innovative protocol that leverages processed hidden representations (§3.1).
  • The security of our protocol is further enhanced by a novel secret-based mechanism (§3.2). We provide a thorough discussion and analysis of various strong and adaptive attack scenarios (§3.3).
@@
  • Our comprehensive experiments with $5$ specified open-source LLMs (from $13$B to $70$B) demonstrate the effectiveness of SVIP: it achieves an average false negative rate of $3.49\%$, while keeping the false positive rate below $3\%$ across $6$ smaller alternative models (§4.1). SVIP introduces negligible overhead (less than $0.01$ seconds per prompt query) for both users and computing providers (§4.2). Furthermore, SVIP can effectively and securely handle $80$ to $120$ million prompt queries in total after a single round of protocol training, with the update mechanism further bolstering security (§4.3).
Table 1: Comparison of simple approaches and our proposed protocol based on the five criteria. A checkmark (✓) indicates that the criterion is satisfied, while a cross (✗) indicates it is not. SVIP is the only method that satisfies all five criteria.
Approach Low FNR Low FPR Efficiency Completion Quality Robustness
Benchmark Prompt Testing
Binary Classifier on Hidden States
SVIP (Ours)
@@ 2 Problem Statement
The verifiable inference problem involves three parties, as illustrated in Figure 1:
@@
    User: An individual who lacks sufficient computing resources and seeks to perform expensive LLM inference tasks on given prompts using decentralized computing providers at a low cost.
@@
    Computing Provider: Decentralized entities, often small companies or individuals, that rent out computational power at competitive prices.
@@
    Platform: A commercial platform that profits by connecting users and computing providers. Importantly, the platform itself does not require significant computational resources, as its primary role is to facilitate and monitor the utilization of computational resources from decentralized providers.
Threat Model
To reduce costs, a computing provider may not actually use the LLM the user specifies. Instead, it may substitute a significantly smaller model, which returns an inferior result. It may also attempt to evade detection by actively concealing dishonest behavior.
The Incentive and Goal of Verifiable Inference
To address this threat, platforms are commercially incentivized to maintain user trust by monitoring provider behavior, ensuring that providers cannot cheat. Trust in the platform underpins its reputation and business model; if users cannot trust that model inference is faithfully executed, they are likely to abandon the platform, leading to significant financial and operational losses.
To mitigate this risk, the platform designs and implements a verification protocol that allows users to verify, with high confidence, whether the computing provider used the specified model for inference. Note that during deployment, the protocol should operate primarily between the user and the provider, with minimal platform involvement. A satisfactory protocol should meet the following criteria: (1) Low False Negative Rate (FNR): The protocol should minimize cases where the computing provider did use the specified LLM for inference but is incorrectly flagged as not using it. (2) Low False Positive Rate (FPR): The protocol should rarely confirm that the computing provider used the specified LLM if it actually used another model. (3) Efficiency: The verification protocol should be computationally efficient and introduce minimal overhead for both the computing provider and the user. (4) Preservation of Completion Quality: The protocol should not compromise the quality of the prompt completion returned by the computing provider. (5) Robustness: The protocol should maintain low FNR and FPR even against adversarial providers attempting to evade detection.
2.1 Simple Approaches Do Not Meet All the Criteria
Table 1 evaluates several straightforward approaches for verifiable LLM inference, as well as our proposed protocol (§3), against the five criteria. All naive approaches fail to satisfy at least one criterion, underscoring the necessity of our method. We exclude solutions that involve multiple computing providers (e.g., cross-verifying results across providers) because such approaches significantly increase user costs, making them impractical for widespread adoption.
Benchmark Prompt Testing
The user curates a small set of prompt examples from established benchmarks and sends them to the computing provider. If the provider’s performance significantly deviates from the reported benchmark metrics for the specified model, the user may suspect dishonest behavior. However, a malicious provider can easily bypass this method by detecting known benchmark prompts and selectively applying the correct model only for those cases, while using an alternative model for all other prompt queries. Additionally, testing such benchmark prompts also increases the user’s inference costs.
Binary Classifier on Hidden States
The user can request the computing provider to return hidden representations from the LLM used for inference, and train a binary classifier on these representations to verify if they come from the specified model. However, a simple attack involves the provider caching hidden representations from the correct model that are unrelated to the user’s input. The dishonest provider could then use a smaller LLM for inference and return these cached irrelevant representations to deceive the classifier while saving costs.
@@ 3 Methodology
Figure 2: Illustration of the motivation behind our framework. The proxy task is trained solely on hidden states from the specified LLM $M_{spec}$. During deployment, strong performance on the proxy task indicates that the provider used $M_{spec}$ as specified, while poor performance suggests otherwise.
@@
It is often challenging to verify whether a computing provider is using an alternative LLM for inference based solely on the returned completion text [^1] To empirically demonstrate this, we train a binary classifier to distinguish between output texts from a specified model (LlaMA-2–13B) and six smaller alternatives. Using 90,000 prompts for training and 10,000 for testing, the classifier (BERT-base-uncased) achieves an FNR of 36.1% and an FPR of 58.9%.. Our framework addresses this by requiring the computing provider to return not only the generated text but also the processed hidden state representations from the LLM inference process.
We design and train a proxy task specifically to perform well only on the hidden representations generated by the specified model during the protocol’s training stage. The intuition behind is that the proxy task transforms the hidden representations into a unique identifier for the model. During deployment, the user can evaluate the performance of the returned hidden states on the proxy task. Strong performance on the proxy task indicates that the correct model was used for inference, while poor performance suggests otherwise. Figure 2 provides an illustration.
Our approach does not depend on expensive cryptographic proofs or protocols, and is highly efficient. Furthermore, it does not involve retraining or fine-tuning the LLMs, operates independently of the model publisher, and can be applied to any LLM with publicly available weight parameters, making it widely applicable.
Notations
Let $x\in\mathcal{V}^{*}$ denote the prompt query from user, where $\mathcal{V}^{*}$ represents the set of all possible string sequences for a vocabulary set $\mathcal{V}$. The specified LLM and alternative LLM are denoted as $\mathcal{M}_{spec}$ and $\mathcal{M}_{alt}$, respectively.
3.1 A Simple Protocol Based on Hidden States
@@
For any LLM $\mathcal{M}$, let $h_{\mathcal{M}}(x)\in\mathbb{R}^{L\times d_{\mathcal{M}}}$ represent the last-layer hidden representations of $x$ produced by $\mathcal{M}$, where $L$ is the length of the tokenized input $x$, and $d_{\mathcal{M}}$ denotes the hidden dimension of $\mathcal{M}$. The computing provider receives $x$ from the user, runs $\mathcal{M}$, and returns $h_{\mathcal{M}}(x)$ to user for subsequent verification. However, to reduce the size of the hidden states returned, we additionally apply a proxy task feature extractor network $g_{\theta}(\cdot):\mathbb{R}^{L\times d_{\mathcal{M}}}\rightarrow\mathbb{R}^{d_{g}}$ parameterized by $\theta$, where $d_{g}$ represents the proxy task feature dimension. The computing provider now also runs $g_{\theta}(\cdot)$ and returns a compressed vector $z(x):=g_{\theta}(h_{\mathcal{M}}(x))$ of dimension $d_{g}$ to the user, significantly reducing the communication overhead. Specifically, for each prompt query, the compressed vector only takes approximately $4$ KB when $d_{g}$ is set to $1024$.
The user is required to perform two tasks locally: obtaining the predicted proxy task output and the label. First, the user runs $f_{\phi}(\cdot)$, using the returned proxy task feature $z(x)$ as input to compute $f_{\phi}(z(x))$. Here, $f_{\phi}(\cdot):\mathbb{R}^{d_{g}}\rightarrow\mathcal{Y}$ is the proxy task head parameterized by $\phi$, where $\mathcal{Y}$ denotes the label space. Second, the user applies a labeling function for the proxy task. We adopt a self-labeling function $y(x):\mathcal{V}^{*}\rightarrow\mathcal{Y}$, which derives the label directly from the input, eliminating the need for external labels or specialized annotators [^2] For instance, we can define $y(x)$ as the Set-of-Words (SoW) representation of the input $x$, which captures the presence of each word in a fixed vocabulary, regardless of frequency. As a concrete example, if $\mathcal{V}=\{a,b,c,d\}$ and $x=``abcc"$, the SoW label $y(x)$ would be a four-dimensional vector $(1,1,1,0)$, indicating whether each token in $\mathcal{V}$ appears in $x$.. The label can either be a scalar or a vector.
@@
With a properly defined loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ and a training dataset $\mathcal{D}$, the platform trains the proxy task according to the following objective:
@@
$\begin{split}\quad&\textbf{Low FNR}:\mathbb{P}\left(V(x,z(x);\phi^{*},\theta^{*})=0~{}\middle|~{}\mathcal{M}_{spec}~{}\text{used}\right)\leq\alpha;\\ \quad&\textbf{Low FPR}:\mathbb{P}\left(V(x,z(x);\phi^{*},\theta^{*})=1~{}\middle|~{}\mathcal{M}_{spec}~{}\text{{not} used}\right)\leq\beta.\end{split}$ (2)
Here $\alpha$ and $\beta$ are predefined thresholds. While a single prompt query may occasionally yield an incorrect verification result due to FNR or FPR, in practice, users can perform the verification over multiple distinct queries and apply a hypothesis testing to reach a final single conclusion with high confidence. Refer to Section 4.1 and Appendix C.2 for a detailed discussion.
(a) Simple Protocol
(b) Secret-based Protocol
Figure 3: Illustration of (a) the simple protocol (Section 3.1); (b) secret-based protocol (Section 3.2).